### v2

@@ 3.5 Characterizations of Efficient Mechanisms
myerson1981optimal proves that (i) an allocation rule can comprise a strategy-proof mechanism if and only if $x$ is monotonic (i.e., $x(v^{\prime})\geq x(v)$ for all $v^{\prime}\geq v$), and (ii) if an allocation is nondecreasing, strategy-proofness uniquely identifies the payment rule (up to constant). Using this approach, Hartline2008 show that the expectation of the residual surplus is equal to the expectation of the virtual valuation for utility $\vartheta$ under strategy-proofness (see their Lemma 2.6):
@@
The problem reduces to the maximization of $\mathbb{E}[\vartheta(v)x(v)]$ subject to the resource constraint, individual rationality, the unit demand condition, and monotonicity of $x$. To do so, we apply a procedure called “ironing” (myerson1981optimal) to obtain a nondecreasing ironed virtual value function $\bar{\vartheta}(v;G_{K})$. Then, to derive an efficient mechanism, starting from higher values of $v$, we greedily increase the allocation probabilities $x(v)$ whereas assigning equal allocation probabilities within intervals where $\bar{\vartheta}(v;G_{K})$ is constant, until the resource constraint $\mathbb{E}[x(v)]\leq\bar{m}$ becomes binding.
@@
The proof proceeds by carefully tracking the mathematical steps involved in the ironing procedure. It is straightforward to see that NBUE is necessary for no screening to be efficient. Under no screening, allocating one good randomly to an agent yields a residual surplus of $\mathbb{E}[v]$ (which is the left-hand side of (14)). In contrast, if we use a take-it-or-leave-it offer requiring a payment $t$, agents accept it if and only if their value $v$ is larger than $t$. By making this offer to $1/\Pr(v>t)$ agents, the planner can allocate the good to one agent on average, generating a residual surplus of $\mathbb{E}[v-t|v>t]$ (which is the right-hand side of (14)). Accordingly, if NBUE is violated, the planner can outperform the no-screening mechanism by (partly) using take-it-or-leave-it offers. Theorem 2 demonstrates not only this necessity but also the sufficiency: when NBUE holds, no mechanism can outperform no screening, for any capacity parameter $\bar{m}$. [^11] As their Corollary 2.11, Hartline2008 provide IHR as a tractable sufficient condition for no screening to be efficient in a finite market. However, since IHR is stronger than NBUE, IHR is not necessary. Alternatively, while IHR is equivalent to a nonincreasing virtual value, this is only a loose sufficient condition for a nonincreasing (constant) ironed virtual value, a property equivalent to NBUE.
@@
The following two statements hold.
@@
    If the reduced value distribution $G_{K}$ has a DHR, then for all $\bar{m}\in(0,1)$, a full-screening mechanism is efficient.
@@
    If the reduced value distribution $G_{K}$ does not have a DHR, then there exists $\bar{m}\in(0,1)$ such that a full-screening mechanism is not efficient.
As previously described, an efficient mechanism is obtained by increasing $x(v)$ from the highest values of $v$. If, at the moment the resource constraint $\mathbb{E}[x(v)]\leq\bar{m}$ becomes binding, the current $v$ does not belong to an interval where the ironed virtual value $\bar{\vartheta}(v;G_{K})$ is flat, then full screening is efficient. Thus, the absence of such intervals—equivalently, a nondecreasing virtual value—is a sufficient condition for full screening efficiency. [^12] For a finite market, Hartline2008 proposes this sufficient condition as Corollary 2.12. This condition corresponds exactly to the distribution $G_{K}$ satisfying DHR. Conversely, if $G_{K}$ does not satisfy DHR, there must exist intervals where the ironed virtual value is flat, implying the existence of some $\bar{m}$ for which full screening is inefficient. Note that full screening may still be efficient even when the ironed virtual value is flat in the low-$v$ region. Thus, DHR is not a necessary condition for full screening to be efficient for every $\bar{m}$.
@@ 3.6 Effects from Increasing Variety
The intuition behind Theorem 4, based on the terminology of survival analysis, is as follows. While $G(v)$ represents the probability that a single product fails by time $v$, its largest order statistic $G_{K}(v)$ represents the probability that all $K$ products fail by time $v$. When calculating the mean residual life of a product, we are guaranteed that all $K$ products are initially functioning. In contrast, when calculating the mean residual life after $t$ periods of use, $\mathbb{E}[v|v>t]$, the condition $v>t$ merely indicates that the largest order statistic $v=\max\{v_{1},\dots,v_{K}\}$ exceeds $t$, ensuring only that at least one product remains functioning at time $t$. Even when conditioning on the event “at least one product survives,” the number of surviving products tends to fall as time $t$ increases. Consequently, when the variety of goods $K$ becomes larger, the mean residual life increases substantially for $t=0$, but this effect is smaller for large $t$. This makes the NBUE condition more likely to hold when we have larger $K$.
@@
Let $r^{\prime}(\cdot;G_{K})$ be the derivative of $r(\cdot;G_{K})$ with respect to $v$. Then, $r^{\prime}(v;G_{K+1})\geq 0$ if $r^{\prime}(v;G_{K})\geq 0$. Furthermore, for all $v\in[0,\bar{v}]$, there exists $K_{0}$ such that for all $K>K_{0}$, $r^{\prime}(v;G_{K})>0$.
@@
Figure 2: The behavior of the distribution of the normalized largest order statistic ($a_{K}v+b_{K}$) when the value distribution is the Weibull distribution with shape parameter $0.9$. The sequence $(a_{K},b_{K})$ is shown in Appendix A.1.
@@ 3.7 Extreme Value Theory
Proposition 2 (fisher1928limiting; Gnedenko1943).
@@
For the proof of Proposition 2, see Theorems 1.1.3 and 1.1.6 of haan2006extreme. [^13] Whereas haan2006extreme employ the representation of generalized extreme value distribution, as they introduce in pp.9-10, it is equivalent to Proposition 2. Theorem 1.1.6 of haan2006extreme additionally implies that, whenever a distribution function $G$ belongs to the domain of attraction of Gumbel, $\hat{G}_{K}(w)\coloneqq G_{K}(a_{K}w+b_{K})\to\Lambda(w)$ with
@@
Reverse-Weibull Uniform, Beta, Reverse-Weibull, etc. (Unimportant for our analysis)
Proposition 2 does not state that the distribution of the normalized maximum converges. [^14] For example, the binomial, Poisson, and geometric distributions do not have an extreme distribution. Nevertheless, it is also known that many continuous distribution functions have an extreme value distribution.151515pickands1975statistical mentions that “Most ‘textbook’ continuous families of distribution functions are discussed in gumbel1958statistics. They all lie in the domain of attraction of some extremal distribution function.” (page 119) Table 1 lists extreme value distributions of common probability distributions. The limit is identical as long as distributions have the same right-tail property, even when their overall shape differs significantly.
@@
The von Mises condition (18) is originally established by VonMises1936 as a tractable sufficient condition for convergence of a distribution function. Later, Pickands1986continuous proves that (18) is a necessary and sufficient condition for convergence of the second derivative.
Proposition 3 (Theorem 5.2 of Pickands1986continuous).
@@ 3.8 Large Market Limit
As the reduced value distribution has neither IHR nor DHR, we apply myerson1981optimal’s (myerson1981optimal) ironing to obtain an efficient mechanism. We can show that there exists a threshold $w^{**}>w^{*}$ such that the efficient mechanism fully screens agents’ preferences for $w>w^{**}$, whereas it conducts no screening for $w<w^{**}$. [^17] The equation (103) in Appendix B characterizes $w^{**}$. If the supply $\bar{m}$ is insufficient to allocate all agents with $w>w^{**}$, then only those with the highest values will win an item. Otherwise, an efficient mechanism first allocates goods with probability one to agents with values $w>w^{**}$. The remaining goods are then distributed equally among the other agents without payment.
@@
When the reduced value distribution function $\hat{G}_{K}$ has an IDHR, there exists $w^{**}>w^{*}$ with which an efficient allocation rule $\hat{x}$ is represented as follows:
@@
    $\hat{x}(w)=\begin{cases}1&\text{ for }w\in[\hat{G}_{K}^{-1}(1-\bar{m}),+\infty),\\ 0&\text{ for }w\in[0,\hat{G}_{K}^{-1}(1-\bar{m})).\end{cases}$ (22)
@@
    $\hat{x}(w)=\begin{cases}1&\text{ for }w\in[w^{**},+\infty),\\ \dfrac{\bar{m}-(1-\hat{G}_{K}(w^{**}))}{\hat{G}_{K}(w^{**})}&\text{ for }w\in[0,w^{**}).\end{cases}$ (23)
(a) Region in which the extreme value distribution has an IHR ($\Phi_{\alpha}(w^{*})$) and the efficient mechanism uses a constant allocation ($\Phi_{\alpha}(w^{**})$)
@@
If $G\in\textbf{VMC}(1/\alpha)$, then for all $\epsilon>0$, there exists $K_{0}\in\mathbb{Z}_{++}$ such that for all $K>K_{0}$, there exists a constant $\bar{x}_{K}\in[0,1]$ such that an efficient allocation rule $\hat{x}$ for $(G,K)$ satisfies $\Pr(\hat{x}(w)=\bar{x}_{K})\in(\Phi_{\alpha}(w^{**}-\epsilon),\Phi_{\alpha}(w^{**}+\epsilon))$, where $w^{**}$ is a solution of the equation (24).
@@ 4 Finite Market
This section considers finite markets, where a finite number of objects are allocated to a finite number of agents (i.e., no continuum copies of agents and objects). The analysis of finite markets is widely believed to be considerably more challenging than analyzing continuous markets. In this study, instead of characterizing efficient mechanisms analytically, we adopted the deep learning-based approach developed by dutting2019optimal, originally for numerically constructing revenue-maximizing auction mechanisms.
@@ 4.2 RegretNet
To numerically derive an efficient mechanism for finite markets, we adopt RegretNet, a deep learning-based method for designing revenue-maximizing auction mechanisms proposed by dutting2019optimal. In RegretNet, a mechanism, which is a function that takes a valuation profile as input and returns an allocation and a payment profile, is represented by a neural network, which can approximate a wide range of functions using a relatively small number of parameters. In dutting2019optimal, the loss function used for learning consists of (i) the negated empirical expected revenue of the seller and (ii) a penalty term for violating strategy-proofness. RegretNet uses the augmented Lagrange method to find a mechanism (i.e., the parameters) minimizing such loss.
@@
As characterized by Hartline2008, SD outperforms VCG if and only if $\alpha>1$, for the single-good case. In the continuous case, SD achieves larger social welfare than VCG as long as the shape parameter is larger than roughly $0.7$, given $K=2$. We observe a similar pattern for finite markets: Even with $c=1$, the threshold at which SD starts to outperform VCG is strictly smaller than $1.0$, and as $c$ increases, the market becomes closer to the continuous case.
@@ 5.1 Problem
The COVID-19 pandemic significantly disrupted global economic and social welfare, prompting rapid vaccine development and deployment. By the end of 2021, around 9.14 billion COVID-19 vaccinations had been administered worldwide (mathieu2021global), marking one of history’s largest allocation problems.
@@ 5.2 Lessons from the COVID-19 Pandemic
During the early rollout of COVID‑19 vaccines, many countries and regions relied on FCFS. These systems turned every available slot into a scarce heterogeneous good and induced high‑intensity competition among potential recipients. In early 2021, several counties in Florida opened public vaccination sites with no reservation system. In Lee County, people paid substantial effort costs (such as camping overnight and queuing for hours) to secure doses until officials acknowledged the chaos and announced a shift to a reservation system (fox_leecounty_2021). Even if a reservation system is in place, similar confusion arises when reservations are accepted on an FCFS basis. In other Florida counties that initially used an FCFS-based reservation system, as our theory predicts, applicants rushed the system immediately after reservations opened and all slots were exhausted within minutes, which in turn led to overwhelmed call centers. Consequently, several local governments quickly abandoned FCFS (routefifty_eventbrite_2021). Similar problems arose in Massachusetts. When the state opened eligibility to older adults, the system experienced delays and outages due to surges in access, leaving many residents unable to make reservations and frustrated (nbcboston_massportal_2021). In Washington, D.C., when an online portal was launched for newly eligible residents, it immediately crashed under heavy traffic, and phone lines were jammed. The incident led to public apologies from officials and a review of the system (dcist_dc_2021). Reservation surges, along with system and call center failures and the resulting public dissatisfaction, were widely observed across many regions in Japan as well (e.g., asahi2021B; asahi2021A; yomiuri2021).
@@
In environments with many heterogeneous reservation slots, our theory implies that such screening-induced races are particularly wasteful, which motivates replacing FCFS with no-screening mechanisms, and several local governments indeed introduced alternative mechanisms. Minnesota introduced a random lottery for seniors over 65. Residents had a 24‑hour window to preregister, and names were drawn at random; those selected received notice that directed them to complete a booking, while the others remained in a queue for future lotteries (startribune_minnesota_2021). Local press reported that the lottery was perceived as fairer and less stressful than the prior FCFS system (postbulletin_minnesota_2021). San Luis Obispo County in California, after facing similar FCFS failures, announced plans to introduce a lottery (mustangnews_slo_2021). A lottery-based mechanism was also introduced in Kakogawa City, Japan, as a direct-mechanism version of random serial dictatorship (Tada2021Kakogawa). There, the high communication cost inherent to direct mechanisms—stemming from the requirement to submit a full preference order—was one reason for later reverting to FCFS.
@@
British Columbia, Canada, introduced a “Get Vaccinated System.” In this reservation system, residents registered once, and the government later sent invitations to book an appointment as supply and age eligibility permitted (bc_getvaccinated_2021). In Singapore, the government created an online pre‑registration form and generated a list of prospective recipients; when vaccine slots became available, those on the list received SMS invitations with booking links (govinsider_singapore_2021). These can be seen as real-world cases where the authorities used a reservation system recommended in this paper.

### v3

@@ 3.5 Characterizations of Efficient Mechanisms
Myerson (1981) proves that (i) an allocation rule can comprise a strategy-proof mechanism if and only if $x$ is monotonic (i.e., $x(v^{\prime})\geq x(v)$ for all $v^{\prime}\geq v$), and (ii) if an allocation is nondecreasing, strategy-proofness uniquely identifies the payment rule (up to constant). Using this approach, Hartline and Roughgarden (2008) show that the expectation of the residual surplus is equal to the expectation of the virtual valuation for utility $\vartheta$ under strategy-proofness (see their Lemma 2.6):
@@
The problem reduces to the maximization of $\mathbb{E}[\vartheta(v)x(v)]$ subject to the resource constraint, individual rationality, the unit demand condition, and monotonicity of $x$. To do so, we apply a procedure called “ironing” (Myerson, 1981) to obtain a nondecreasing ironed virtual value function $\bar{\vartheta}(v;G_{K})$. Then, to derive an efficient mechanism, starting from higher values of $v$, we greedily increase the allocation probabilities $x(v)$ whereas assigning equal allocation probabilities within intervals where $\bar{\vartheta}(v;G_{K})$ is constant, until the resource constraint $\mathbb{E}[x(v)]\leq\bar{m}$ becomes binding.
@@
Note that statement (ii) is stronger than the converse of statement (i), in the sense that the quantifier over $\bar{m}$ is “for all” rather than “there exists.” When the reduced value distribution $G_{K}$ is not NBUE, we can improve residual surplus by screening out agents with sufficiently low values, at least in a neighborhood of $v=0$. As a result, regardless of the value of $\bar{m}\in(0,1)$, a no-screening mechanism cannot be efficient.
The proof proceeds by carefully tracking the mathematical steps involved in the ironing procedure. It is straightforward to see that NBUE is necessary for no screening to be efficient. Under no screening, allocating one good randomly to an agent yields a residual surplus of $\mathbb{E}[v]$ (which is the left-hand side of (14)). In contrast, if we use a take-it-or-leave-it offer requiring a payment $t$, agents accept it if and only if their value $v$ is larger than $t$. By making this offer to $1/\Pr(v>t)$ agents, the planner can allocate the good to one agent on average, generating a residual surplus of $\mathbb{E}[v-t|v>t]$ (which is the right-hand side of (14)). Accordingly, if NBUE is violated, the planner can outperform the no-screening mechanism by (partly) using take-it-or-leave-it offers. Theorem 2 demonstrates not only this necessity but also the sufficiency: when NBUE holds, no mechanism can outperform no screening, for any capacity parameter $\bar{m}$. [^11] As their Corollary 2.11, Hartline and Roughgarden (2008) provide IHR as a tractable sufficient condition for no screening to be efficient in a finite market. However, since IHR is stronger than NBUE, IHR is not necessary. Alternatively, while IHR is equivalent to a nonincreasing virtual value, this is only a loose sufficient condition for a nonincreasing (constant) ironed virtual value, a property equivalent to NBUE.
@@
The following two statements are equivalent.
@@
    The reduced value distribution $G_{K}$ has a DHR.
@@
    For all $\bar{m}\in(0,1)$, a full-screening mechanism is efficient.
As previously described, an efficient mechanism is obtained by increasing $x(v)$ from the highest values of $v$. If, at the moment the resource constraint $\mathbb{E}[x(v)]\leq\bar{m}$ becomes binding, the current $v$ does not belong to an interval where the ironed virtual value $\bar{\vartheta}(v;G_{K})$ is flat, then full screening is efficient. Thus, the absence of such intervals—equivalently, a nondecreasing virtual value—is a sufficient condition for full screening efficiency. [^12] For a finite market, Hartline and Roughgarden (2008) proposes this sufficient condition as Corollary 2.12. This condition corresponds exactly to the distribution $G_{K}$ satisfying DHR. Conversely, if $G_{K}$ does not satisfy DHR, there must exist intervals where the ironed virtual value is flat, implying the existence of some $\bar{m}$ for which full screening is inefficient. Note that full screening may still be efficient even when the ironed virtual value is flat in the low-$v$ region. Thus, DHR is not a necessary condition for full screening to be efficient for some fixed $\bar{m}$.
@@ 3.6 Effects from Increasing Variety
The intuition behind Theorem 4, based on the terminology of survival analysis, is as follows. While $G(v)$ represents the probability that a single product fails by time $v$, its largest order statistic $G_{K}(v)$ represents the probability that all $K$ products fail by time $v$. When calculating the mean residual life of a product, we are guaranteed that all $K$ products are initially functioning. In contrast, when calculating the mean residual life after $t$ periods of use, $\mathbb{E}[v|v>t]$, the condition $v>t$ merely indicates that the largest order statistic $v=\max\{v_{1},\dots,v_{K}\}$ exceeds $t$, ensuring only that at least one product remains functioning at time $t$. Even when conditioning on the event “at least one product survives,” the number of surviving products tends to fall as time $t$ increases. Consequently, when the variety of goods $K$ becomes larger, the mean residual life increases substantially for $t=0$, but this effect is smaller for large $t$. This makes the NBUE condition more likely to hold when we have a larger $K$.
@@
Let $r^{\prime}(\cdot;G_{K})$ be the derivative of $r(\cdot;G_{K})$ with respect to $v$. Then, $r^{\prime}(v;G_{K+1})\geq 0$ if $r^{\prime}(v;G_{K})\geq 0$. Furthermore, for all $v\in(0,\bar{v})$, there exists $K_{0}$ such that for all $K>K_{0}$, $r^{\prime}(v;G_{K})>0$.
@@
Figure 2: The behavior of the distribution of the normalized largest order statistic ($a_{K}w+b_{K}$) when the value distribution is the Weibull distribution with shape parameter $0.9$. The sequence $(a_{K},b_{K})$ is shown in Appendix A.1.
@@ 3.7 Extreme Value Theory
Proposition 2 (Fisher and Tippett (1928); Gnedenko (1943)).
@@
For the proof of Proposition 2, see Theorems 1.1.3 and 1.1.6 of De Haan and Ferreira (2006). [^13] Whereas De Haan and Ferreira (2006) employ the representation of generalized extreme value distribution, as they introduce in pp.9-10, it is equivalent to Proposition 2. Theorem 1.1.6 of De Haan and Ferreira (2006) additionally implies that, whenever a distribution function $G$ belongs to the domain of attraction of Gumbel, $\hat{G}_{K}(w)\coloneqq G_{K}(a_{K}w+b_{K})\to\Lambda(w)$ with
@@
Reverse-Weibull Uniform, Beta, Reverse-Weibull, etc.
Proposition 2 does not state that the distribution of the normalized maximum converges. [^14] For example, the binomial, Poisson, and geometric distributions do not have an extreme distribution. Nevertheless, it is also known that many continuous distribution functions have an extreme value distribution. [^15] Pickands III (1975) mentions that “Most ‘textbook’ continuous families of distribution functions are discussed in Gumbel (1958). They all lie in the domain of attraction of some extremal distribution function.” (page 119) Table 1 lists extreme value distributions of common probability distributions. The limit is identical as long as distributions have the same right-tail property, even when their overall shape differs significantly.
@@
The von Mises condition (18) is originally established by von Mises (1936) as a tractable sufficient condition for convergence of a distribution function. Later, Pickands III (1986) proves that (18) is a necessary and sufficient condition for convergence of the second derivative.
Proposition 3 (Theorem 5.2 of Pickands III (1986)).
@@ 3.8 Large Market Limit
As the reduced value distribution has neither IHR nor DHR, we apply Myerson’s (1981) ironing to obtain an efficient mechanism. Under IDHR, the ironing yields a cutoff $w^{**}\in(w^{*},+\infty]$ such that the mechanism conducts full screening only for types above $w^{**}$. If $w^{**}<+\infty$, the mechanism screens agents with $w>w^{**}$, whereas it does not screen those with $w<w^{**}$. Note that a distribution can be NBUE even when it has a DHR, and in such a case, $w^{**}=+\infty$ and a no screening mechanism is efficient, whereas Fréchet is not NBUE. [^17] If $w^{**}<+\infty$, the threshold $w^{**}$ is characterized by (103) in Appendix B.7. The mean residual life function $\mathbb{E}[v-t|v>t]$ of Fréchet diverges as $t$ increases, and thus it is not NBUE. See Section 6.2 of Embrechts et al. (2013), for example. If the supply $\bar{m}$ is insufficient to allocate all agents with $w>w^{**}$, then only those with the highest values will win an item. Otherwise, an efficient mechanism first allocates goods with probability one to agents with values $w>w^{**}$. The remaining goods are then distributed equally among the other agents without payment.
@@
When the reduced value distribution function $\hat{G}_{K}$ has an IDHR, there exists $w^{**}\in(w^{*},+\infty]$ with which an efficient allocation rule $\hat{x}$ is represented as follows:
@@
    $\hat{x}(w)=\begin{cases}1&\text{ for }w\in[\hat{G}_{K}^{-1}(1-\bar{m}),+\infty),\\ 0&\text{ for }w\in(-\infty,\hat{G}_{K}^{-1}(1-\bar{m})).\end{cases}$ (22)
@@
    $\hat{x}(w)=\begin{cases}1&\text{ for }w\in[w^{**},+\infty),\\ \dfrac{\bar{m}-(1-\hat{G}_{K}(w^{**}))}{\hat{G}_{K}(w^{**})}&\text{ for }w\in(-\infty,w^{**}).\end{cases}$ (23)
In particular, if $w^{**}=+\infty$, then (ii) implies $\hat{x}(w)\equiv\bar{m}$ (i.e., no screening).
(a) Cumulative probabilities of regions in which the extreme value distribution has an IHR ($\Phi_{\alpha}(w^{*})$) and the efficient mechanism uses a constant allocation ($\Phi_{\alpha}(w^{**})$)
@@
If $G\in\textbf{VMC}(1/\alpha)$, then for all $\epsilon>0$, there exists $K_{0}\in\mathbb{Z}_{++}$ such that for all $K>K_{0}$, there exists a constant $\bar{x}_{K}\in[0,1]$ such that an efficient allocation rule $\hat{x}$ for $(G,K)$ satisfies $\hat{x}(w)=\bar{x}_{K}$ for all $w\in(\epsilon,w^{**}-\epsilon)$, where $w^{**}$ is a solution of the equation (24).
@@ 4 Finite Market
This section considers finite markets, where a finite number of objects are allocated to a finite number of agents (i.e., no continuum copies of agents and objects). The analysis of finite markets is widely believed to be considerably more challenging than analyzing continuous markets. In this study, instead of characterizing efficient mechanisms analytically, we adopted the deep learning-based approach developed by Dütting et al. (2019), originally for numerically constructing revenue-maximizing auction mechanisms.
@@ 4.2 RegretNet
To numerically derive an efficient mechanism for finite markets, we adopt RegretNet, a deep learning-based method for designing revenue-maximizing auction mechanisms proposed by Dütting et al. (2019). In RegretNet, a mechanism, which is a function that takes a valuation profile as input and returns an allocation and a payment profile, is represented by a neural network, which can approximate a wide range of functions using a relatively small number of parameters. In Dütting et al. (2019), the loss function used for learning consists of (i) the negated empirical expected revenue of the seller and (ii) a penalty term for violating strategy-proofness. RegretNet uses the augmented Lagrange method to find a mechanism (i.e., the parameters) minimizing such loss.
@@
As characterized by Hartline and Roughgarden (2008), SD outperforms VCG if and only if $\alpha>1$, for the single-good case. In the continuous case, SD achieves larger social welfare than VCG as long as the shape parameter is larger than roughly $0.7$, given $K=2$. We observe a similar pattern for finite markets: Even with $c=1$, the threshold at which SD starts to outperform VCG is strictly smaller than $1.0$, and as $c$ increases, the market becomes closer to the continuous case.
@@ 5.1 Problem
The COVID-19 pandemic significantly disrupted global economic and social welfare, prompting rapid vaccine development and deployment. By the end of 2021, around 9.14 billion COVID-19 vaccinations had been administered worldwide (Mathieu et al., 2021), marking one of history’s largest allocation problems.
@@ 5.2 Lessons from the COVID-19 Pandemic
During the early rollout of COVID‑19 vaccines, many countries and regions relied on FCFS. These systems turned every available slot into a scarce heterogeneous good and induced high‑intensity competition among potential recipients. In early 2021, several counties in Florida opened public vaccination sites with no reservation system. In Lee County, people paid substantial effort costs (such as camping overnight and queuing for hours) to secure doses until officials acknowledged the chaos and announced a shift to a reservation system (Fox 4 News, 2021). Even if a reservation system is in place, similar confusion arises when reservations are accepted on an FCFS basis. In other Florida counties that initially used an FCFS-based reservation system, as our theory predicts, applicants rushed the system immediately after reservations opened and all slots were exhausted within minutes, which in turn led to overwhelmed call centers. Consequently, several local governments quickly abandoned FCFS (Coleman, 2021). Similar problems arose in Massachusetts. When the state opened eligibility to older adults, the system experienced delays and outages due to surges in access, leaving many residents unable to make reservations and frustrated (NBC Boston, 2021). In Washington, D.C., when an online portal was launched for newly eligible residents, it immediately crashed under heavy traffic, and phone lines were jammed. The incident led to public apologies from officials and a review of the system (Barthel et al., 2021). Reservation surges, along with system and call center failures and the resulting public dissatisfaction, were widely observed across many regions in Japan as well (e.g., Asahi Shimbun, 2021a, b; Yomiuri Shimbun, 2021).
@@
In environments with many heterogeneous reservation slots, our theory implies that such screening-induced races are particularly wasteful, which motivates replacing FCFS with no-screening mechanisms, and several local governments indeed introduced alternative mechanisms. Minnesota introduced a random lottery for seniors over 65. Residents had a 24‑hour window to preregister, and names were drawn at random; those selected received notice that directed them to complete a booking, while the others remained in a queue for future lotteries (Olson, 2021). Local press reported that the lottery was perceived as fairer and less stressful than the prior FCFS system (Stolle, 2021). San Luis Obispo County in California, after facing similar FCFS failures, announced plans to introduce a lottery (Pitcher, 2021). A lottery-based mechanism was also introduced in Kakogawa City, Japan, as a direct-mechanism version of random serial dictatorship (Tada, 2021). There, the high communication cost inherent to direct mechanisms—stemming from the requirement to submit a full preference order—was one reason for later reverting to FCFS.
@@
British Columbia, Canada, introduced a “Get Vaccinated System.” In this reservation system, residents registered once, and the government later sent invitations to book an appointment as supply and age eligibility permitted (Government of British Columbia, 2021). In Singapore, the government created an online pre‑registration form and generated a list of prospective recipients; when vaccine slots became available, those on the list received SMS invitations with booking links (Liew, 2022). These can be seen as real-world cases where the authorities used a reservation system recommended in this paper.